# Enhanced quantum radiation with flying-focus laser pulses

### v1

@@ Abstract
The saturation of the electron energy loss at higher pulse energies suggests that an observable without a theoretical upper limit would better elucidate the benefits of FF pulses in these conditions. One such observable, that is also important for applications, is the probability $P$ of photon emission per unit time. Multiplying the probability by the number of electrons colliding with the laser pulse provides the average number of emitted photons. This number is exact for uncorrelated photon emission [52, 53] and is an excellent approximation when the contribution from correlated photon emission is negligible, as is the case here (see Section I of the SM [58]).
The interpolation formula for the probability of single photon emission per unit time within the LCFA is given by [50]
@@
which is better than 2% accurate for all values of $\chi$. In the classical regime ($\chi\ll 1$), $dP/dt\propto\chi/\mathcal{E}$; while in the quantum regime ($\chi\gg 1$), $dP/dt\propto\chi^{2/3}/\mathcal{E}$. These proportionalities indicate that the emission probability $P$ scales more favorably with the interaction time than the laser intensity in both regimes. Thus, for fixed collision parameters (i.e., pulse energy, electron energy, focusing), FF pulses will have an advantage over SFG pulses regardless of $\chi$.
@@
For definiteness, consider the most favorable conditions for a SFG pulse, which is the classical regime where $\chi\ll 1$. Maximizing the emission probability for a SFG pulse requires matching the interaction time $t_{\text{int}}$ to the Rayleigh range $z_{R}$, so that electrons spend much of the interaction near the focus where the intensity is high. The optimum is $t_{\text{int}}\approx 6z_{R}$, which yields $P_{\text{C,SFG}}\propto U^{1/2}$ (see Section II of SM [58] for derivation). With a FF pulse, the focus moves with the electrons, so the entire interaction time is spent near the focus where the intensity is high. As a result, the interaction time can be independent of the FF Rayleigh range $z_{R,\text{FF}}\equiv\omega_{0}\sigma_{0}^{2}$. In this case, $P_{\text{C,FF}}\propto(t_{\text{int}}/z_{R,\text{FF}})^{1/2}U^{1/2}$, and the optimum is achieved by making the interaction time as long as possible. The ratio $P_{\text{C,\text{FF}}}/P_{\text{C,SFG}}\propto(t_{\text{int}}/z_{R,\text{FF}})^{1/2}$ reveals that, for fixed laser pulse energy, a FF pulse will have an advantage over a SFG pulse when the interaction time of the FF is longer than its Rayleigh range.
The advantage of FF pulses is confirmed in Fig. 3, which compares the number of photons $\hat{N}_{\gamma}$ emitted per unit photon energy per electron from the collision of $\mathcal{E}_{0}=$ 10 GeV electrons with $U=$ 10 J pulses. The parameters were chosen to be experimentally feasible [6]: the wavelength and spot size of the laser pulses were $\lambda_{0}=1\ \mu$m and $\sigma_{0}=2\ \mu$m, respectively, and the electron beam was initialized with normally distributed transverse positions with a variance of $0.75\ \mu$m, a divergence of 1 mrad, and a length of $5\lambda_{0}$. The remaining parameters were determined by maximizing the photon yields from the SFG and FF pulses while still satisfying the LCFA: The FF pulse (solid line) had $\xi_{0}=6$ and $t_{\text{int}}=1.61$ ps, and the SFG pulse (dashed line) had $\xi_{0}=16$ and $t_{\text{int}}=227$ fs. Note that the choice of $\xi_{0}=6$ in the FF case (as opposed to $\xi_{0}=5$ used in Fig. 1) ensures the validity of the LCFA for off-axis electrons that experience lower field strengths. The photon spectra were generated using a Monte Carlo code (see Section I of SM [58]).
@@
Figure 3 shows that the FF pulse results in more photons per electron than the SFG pulse by a factor of 7 in the 1 - 2 MeV range, a factor of 5 in the 2 - 10 MeV, and a factor of 4 in 10 - 20 MeV range. Based on these results, the collision of a 100 pC beam of $\mathcal{E}_{0}$ = 10 GeV electrons with a FF pulse could produce $5\times 10^{10}$ photons in the 1 - 20 MeV range, compared to only $9\times 10^{9}$ photons with a SFG pulse. Moreover, despite its much lower amplitude ($\xi_{0}=6$), the FF pulse still produces a substantial number of high-energy photons ($\omega\sim\mathcal{E}_{0}$) nearly equal to that produced by the higher amplitude ($\xi_{0}=16$) SFG pulse.
@@
Figure 3: Spectra of emitted photons from the head-on collision of $\mathcal{E}_{0}=10$ GeV electrons with $U=10$ J laser pulses. The inset shows the number of photons emitted as a function of energy and $\eta_{\text{LCFA}}^{2/3}$. The number of photons emitted per electron in each shaded interval is displayed in the bottom left. The FF pulse (blue solid) results in approximately 5×5$\times$5 × more photons in the 1-20 MeV range than the SFG pulse (orange dashed).
@@
Because the advantage of the FF pulse occurs in the low-energy part of the spectrum, it is important to verify that the breakdown of the LCFA at low photon energies does not invalidate the results [see Eq. (5) and Refs. [59, 60, 61]]. To assess possible deviations from the LCFA in this part of the spectrum, simulations were performed with the particle-in-cell (PIC) code Smilei [63] including the SFQED-tookit [64], which implements corrections beyond the LCFA [60]. These simulations are in excellent agreement with the Monte Carlo results presented in Fig. 3 (see Section III of SM [58]). As additional verification, the inset of Fig. 3 shows that almost all photons in the 1 - 20 MeV range are emitted with $\eta_{\text{LCFA}}^{2/3}<0.1$.
@@
Finally, the Monte Carlo code used to generate the photon spectra in Fig. 3 does not account for the decay of the emitted photons into electron-positron pairs, which would also radiate [65]. As demonstrated in Section IV of the SM [58], pair production can be safely neglected for both the SFG and FF pulses considered in Fig. 3. At higher field strengths ($\xi_{0}\gtrsim 40$), pair production in a SFG pulse would deplete the most energetic photons and slightly increase the low-energy photon yield, but the overall energy loss and photon yield would be suppressed (see Section III of the SM [58]).
In conclusion, we have shown that a flying focus (FF) pulse can significantly enhance the signatures of quantum radiation reaction when compared to a stationary focus Gaussian (SFG) pulse with the same energy. First, an electron can radiate significantly more energy in a FF pulse than in a SFG pulse. Second, and perhaps more importantly, a FF pulse enhances the yield of emitted photons, especially in the low-energy part of the spectrum. As an example, a 10 GeV electron colliding with a 10 J flying focus pulse produces 5×5$\times$5 × more 1 - 20 MeV photons than with a SFG pulse. From a practical standpoint, this enhancement could benefit applications, such as radiosurgery, nuclear chemistry, and non-destructive imaging.
Acknowledgements.
@@
The authors thank Samuele Montefiori and Matteo Tamburini for kindly sharing a development version of their SFQED-toolkit module [64] for SMILEI and providing support. This project has received funding from the European Union’s Horizon Europe research and innovation program under the Marie Skłodowska-Curie grant agreement No. 101105246-STEFF. This material is based upon work supported by the U.S. Department of Energy [National Nuclear Security Administration] University of Rochester “National Inertial Confinement Fusion Program” under Award Number DE-NA0004144 and U.S. Department of Energy, Office of Science, under Award Number DE-SC0021057. This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.

### v2

@@ II.1 Photon yield
The saturation of the electron energy loss at higher pulse energies suggests that an observable without a theoretical upper limit would better elucidate the benefits of FF pulses in these conditions. One such observable, that is also important for applications, is the total number of emitted photons. Before presenting numerical calculations of the photon number, it is instructive to examine its scaling behavior to explain the advantages of a FF pulse compared to a SFG pulse. Within the LFCA, the probability $P$ of single photon emission per unit time can be calculated using the interpolation formula [Baier:1998vh]
@@
which is better than 2% accurate for all values of $\chi$. Multiplying the time-integrated probability by the number of electrons colliding with the laser pulse provides an estimate of the average number of emitted photons. Provided that the energy loss is accounted for as discussed below Eq. (1), this estimate is quantitatively accurate if higher-order processes, such as multiple photon emission within a formation length can be neglected.
@@
In the classical regime ($\chi\ll 1$), Eq. (6) scales as $dP/dt\propto\chi/\mathcal{E}$; while in the quantum regime ($\chi\gg 1$), $dP/dt\propto\chi^{2/3}/\mathcal{E}$. These proportionalities indicate that the emission probability $P$ scales more favorably with the interaction time than the laser intensity in both regimes. Thus, for fixed collision parameters (i.e., pulse energy, electron energy, focusing), FF pulses have an advantage over SFG pulses regardless of $\chi$.
@@
To demonstrate this, we first consider the most favorable conditions for a SFG pulse, which is the classical regime where $\chi\ll 1$. Maximizing the emission probability for a SFG pulse requires matching the interaction time $t_{\text{int}}$ to the Rayleigh range $z_{R}$, so that electrons spend much of the interaction near the focus where the intensity is high. The optimum is $t_{\text{int}}\approx 5z_{R}-7z_{R}$, which yields $P_{\text{C,SFG}}\propto U^{1/2}$ (see Section II of SM [supplemental]). In a FF pulse, the focus moves with the electrons, so the entire interaction time is spent near the focus. As a result, the interaction time is independent of the FF Rayleigh range $z_{R,\text{FF}}\equiv\omega_{0}\sigma_{0}^{2}$. In this case, $P_{\text{C,FF}}\propto(t_{\text{int}}/z_{R,\text{FF}})^{1/2}U^{1/2}$, such that the average number of emitted photons increases with the interaction time. The ratio $P_{\text{C,\text{FF}}}/P_{\text{C,SFG}}\propto(t_{\text{int}}/z_{R,\text{FF}})^{1/2}$ reveals that, for fixed laser pulse energy, a FF pulse will have an advantage over a SFG pulse when the interaction time of the FF is longer than its Rayleigh range. In the deep quantum regime ($\chi\gg 1$) the scaling is even more favorable, starting as $P_{\text{Q,FF}}\propto U^{1/3}t_{\text{int}}^{2/3}$ before asymptoting to the classical scaling as the electron loses energy.
Quantum enhancement.
Figure 3 demonstrates that a FF pulse can yield more photons emitted in the quantum regime than SFG pulses with the same energy. In the on-axis collision of 10 GeV electrons with a 1 J, $\xi_{0}=6$ FF pulse (as considered in Fig. 1a), 10.0 photons were emitted at $\chi>0.1$ per electron in the energy range 100 MeV – 2 GeV compared to 5.2 photons per electron with a 1 J, $\xi_{0}=13$ SFG pulse (maximum energy loss from Fig. 1a) and compared to 6.3 photons per electron in a 1 J, $\xi_{0}=9$ SFG pulse. The amplitude, spot size, and interaction time ($t_{\text{int}}=128$ fs) of the $\xi_{0}=9$ SFG pulse were optimized to achieve the highest photon yield with $\chi>0.1$ in the 100 MeV – 2 GeV range. The electrons lost 84%, 64%, and 59% of their initial energy in the FF, $\xi_{0}=13$ SFG, and $\xi_{0}=9$ SFG pulses, respectively. The electron energy distributions had initial standard deviation widths of 1%, which were broadened by quantum photon emission to 39%, 50%, and 38% with maximum quantum parameters $\chi_{0}\approx 0.59$, $1.23$, and $0.83$ [Neitz_2013, Neitz_2014, Vranic:2015sft]. The photon spectra were generated using the Monte Carlo code described in Section I of SM [supplemental]. Photons emitted with energies below 100 MeV in the FF collision are not reported because they do not meet the requirement for satisfying the LCFA approximation discussed below Eq.(5).
Photon yield increase.
[…]
@@
Figure 4 shows that the FF pulse colliding with 10 GeV electrons results in more photons per electron than the SFG pulse by a factor of 7 in the 1 – 2 MeV range, a factor of 5 in the 2 – 10 MeV, and a factor of 4 in 10 – 20 MeV range. Based on these results, the collision of a 100 pC beam of $\mathcal{E}_{0}$ = 10 GeV electrons with a FF pulse could produce $5\times 10^{10}$ photons in the 1 – 20 MeV range, compared to only $9\times 10^{9}$ photons with a SFG pulse.
Despite its much lower amplitude ($\xi_{0}=6$), the FF pulse produces a substantial number of high-energy photons ($\omega_{0}>20$ MeV) comparable to that produced by the higher amplitude ($\xi_{0}=16$) SFG pulse. However, most of the photons are in the MeV range. For 10 GeV electrons, there are two-times more photons in the 1–20 MeV range (accounting for 5% of $\mathcal{E}_{0}$ on average) than above 20 MeV (accounting for 92% of $\mathcal{E}_{0}$). The maximum quantum parameter in the FF collision is $\chi_{0}\approx 0.6$ and many photons are emitted in the quantum regime of radiation reaction.
For a FF pulse colliding with 1 GeV electrons, there are eight-times more photons in the 1–20 MeV range than above 20 MeV. The 1–20 MeV photons account for 36% of $\mathcal{E}_{0}$ on average, while the photons above 20 MeV account for 34%. The higher photon yield produced by the FF is due to a larger number of photons below 20 MeV. As $\chi_{0}\approx 0.06$, the classical scaling applies, and the 3 times increase in the number of 1 – 20 MeV photons is comparable to the $\sim$2.7 times overall increase predicted by the classical scaling, $dP/dt\propto\chi/\mathcal{E}$.
@@
Figure 4: Spectra of emitted photons from the head-on collision of $\mathcal{E}_{0}=10$ GeV and 1 GeV electrons with $U=10$ J laser pulses. The inset shows the number of photons emitted as a function of energy and $\eta_{\text{LCFA}}^{2/3}$ for the FF pulse collisions. The number of photons emitted per electron in each shaded interval is displayed in the bottom left. The FF pulse (solid lines) results in approximately five-times and three-times more photons in the 1–20 MeV range than the SFG pulse (dashed lines) for 10 and 1 GeV electrons, respectively.
III Simulation validation
@@
Because the advantage of the FF pulse occurs in the low-energy part of the spectrum, it is important to verify that the breakdown of the LCFA at low photon energies does not invalidate the results [see Eq. (5) and Refs. [Di_Piazza_2018, Di_Piazza_2019, Ilderton_2019_b]]. To assess possible deviations from the LCFA in this part of the spectrum, simulations were performed with the particle-in-cell (PIC) code Smilei [DEROUILLAT2018351] including the SFQED-tookit [Montefiori:2023mkc], which implements corrections beyond the LCFA [Di_Piazza_2019]. These simulations are in excellent agreement with the Monte Carlo results presented in Figs. 3 and 4 (see Section III of SM [supplemental]). As additional verification, the insets of Figs. 3 and 4 show that almost all photons in the studied ranges are emitted with $\eta_{\text{LCFA}}^{2/3}<0.1$.
@@
Finally, the Monte Carlo code used to generate the photon spectra in Figs. 3 and 4 does not account for the decay of the emitted photons into electron-positron pairs, which would also radiate [mercuri2021impact]. As demonstrated in Section IV of the SM [supplemental], pair production can be safely neglected for both the SFG and FF pulses considered in Figs. 3 and 4. At higher field strengths ($\xi_{0}\gtrsim 40$), pair production in a SFG pulse would deplete the most energetic photons and slightly increase the low-energy photon yield, but the overall energy loss and photon yield would be suppressed (see Section III of the SM [supplemental]).
IV Conclusions
In conclusion, we have shown that a flying focus (FF) pulse can significantly enhance the signatures of quantum radiation reaction when compared to a stationary focus Gaussian (SFG) pulse with the same energy. First, an electron can radiate significantly more energy in a FF pulse than in a SFG pulse. Second, a FF pulse enhances the yield of emitted photons, especially in the low-energy part of the spectrum. Moreover, the number of photons emitted in the quantum regime of the interaction (with $\chi>0.1$) can be larger when using a FF in place of a SFG pulse. These improvements are a direct result of the energy loss in the quantum regime and the photon emission probability in both the quantum and classical regimes scaling more favorably with the interaction time than the field intensity. Thus, any Compton scattering configuration can use a FF pulse in place of a SFG pulse to exploit this advantage. Furthermore, the use of FF pulses lowers the required power and intensity, bypasses the need for larger compression gratings, and provides a more controllable environment with simpler diagnostics of the interaction field strength [Popruzhenko:2023tqx], while simultaneously enhancing the radiation signatures.
@@ IV Conclusions
Acknowledgments. The authors thank Samuele Montefiori and Matteo Tamburini for kindly sharing a development version of their SFQED-toolkit module [Montefiori:2023mkc] for SMILEI and providing support. We also thank Elias Gerstmayr for discussions regarding laboratory sources of gamma rays. This project has received funding from the European Union’s Horizon Europe research and innovation program under the Marie Skłodowska-Curie grant agreement No. 101105246-STEFF. This work was supported by the Ministry of Education, Youth and Sports of the Czech Republic through the e-INFRA CZ (ID:90254). This material is based upon work supported by the U.S. Department of Energy [National Nuclear Security Administration] University of Rochester “National Inertial Confinement Fusion Program” under Award Number DE-NA0004144 and U.S. Department of Energy, Office of Science, under Award Number DE-SC0021057.
This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.
See pages 1 of supplementalSee pages 2 of supplementalSee pages 3 of supplementalSee pages 4 of supplementalSee pages 5 of supplementalSee pages 6 of supplemental